# Differential privacy guarantees of Markov chain Monte Carlo algorithms

### v1

@@ 1 Introduction
Several works obtain DP of the one-step transitions of the MCMC algorithm by injecting extra noise in various ways, e.g. in the acceptance-rejection step of Metropolis-Hastings-type algorithms (Heikkilä et al., 2019; Yıldırım & Ermics, 2019), or leveraging the additional randomness already introduced by subsampling strategies (Wang et al., 2015; Bierkens & Duncan, 2022), optionally together with clipping of the gradient that drives the moves of the chain (Song et al., 2013; Abadi et al., 2016). Many iterations of the MCMC chain are required to approach the posterior distribution, and the typical strategy to obtain guarantees on the DP of the algorithm after some number $n$ of iterations is to use composition bounds (Abadi et al., 2016; Wang et al., 2019; Bassily et al., 2014). However, the privacy implied using composition bounds for Markov chains naturally decays with the number of steps (Kairouz et al., 2015) and as a consequence the injected noise must scale with the number of iterations in order to obtain uniform privacy bounds for the final draw. This limitation was recently addressed by e.g.,Chourasia et al. (2021) and Altschuler & Talwar (2022), who obtained uniform in time bounds on the (Rényi) DP of gradient descent-type algorithms. However, these results are limited to the strongly convex case or have bad dependence on the step-size and diameter of the state space.
@@
In the first part (Section 3) we consider general MCMC algorithms and establish a clear connection between the DP guarantees of the posterior distribution and of a corresponding MCMC algorithm. In particular, we find that the DP of the posterior is the crucial starting point to perform Bayesian inference. This fact is supported by Proposition 3.4, which shows that if the posterior has weaker DP guarantees than the MCMC algorithm, the law of the MCMC chain after $n$ iterations is far from the posterior in total variation distance. This result emphasises the importance of starting with a differentially private model, a property that has been mostly neglected in the literature. The other results presented in Section 3 (Propositions 3.2 and 3.10) assume DP of the posterior and use it to show $(\varepsilon,\delta)$-DP for the MCMC algorithm when we either release the $n$-th state of the chain or a (noisy) Monte Carlo estimator. These bounds are tight when the algorithm is close to equilibrium.
In Section 4, we introduce a novel approach to establish $(\varepsilon,\delta)$-DP and Rényi-DP of the full path of an MCMC algorithm, or also only of its state after $n$ iterations, in a non-convex setting. Our approach, presented in Section 4.1, is based on obtaining high probability bounds of the Radon-Nikodym derivative of the law of the output obtained for two different, adjacent datasets. In Section 4.2 we illustrate how this strategy can be applied to the case of MCMC algorithms based on diffusions, where we rely on Girsanov’s theorem to express the Radon-Nikodym derivative. Our strategy of proof allows to obtain bounds for the entire trajectory of a variety of stochastic algorithms, significantly improving on known composition bounds. In particular, we show that the entire trajectory up to the $n$-th iteration is $(\varepsilon,\delta)$-DP, where $\varepsilon$ is $O(n+\sqrt{n\log(1/\delta)})$ and $\delta>0$ is constant, and also $(\alpha,\varepsilon)$-Rényi-DP with $\varepsilon=O(\alpha n)$ for $\alpha>0$ constant. Thanks to a careful perturbation technique, we are also able to apply our approach to the case where only the final state after $n$ iterations is released, and not the entire path. In this case, we obtain uniform in time guarantees both for DP and Rényi-DP. We then specialise our analysis to two classical MCMC algorithms based on the Langevin diffusion: the unadjusted Langevin algorithm and its stochastic gradient variant (see Section 5). Leveraging our approach, we address the open problem of obtaining uniform in time DP guarantees in a non-convex setting.
@@ 2 Preliminaries
$\pi_{\mathcal{D}}(\mathsf{B})=\left.\int_{\mathsf{B}}L_{\mathcal{D}}(\theta)\eta(\mathrm{d}\theta)\middle/{Z_{\mathcal{D}}}\right.\;,\quad\mathsf{B}\in\mathcal{B}(\mathrm{E})\;,$ (1)
where $(\mathrm{E},\mathcal{B}(\mathrm{E}))$ is a Borel space, $L_{\mathcal{D}}(\theta)$ is the likelihood associated with ${\mathcal{D}}$, $\eta$ is a prior distribution for the parameter, and $Z_{\mathcal{D}}=\int_{\mathrm{E}}L_{\mathcal{D}}(\theta)\eta(\mathrm{d}\theta)$. In order to make practical use of the Bayesian framework, it is then essential to have access to key statistics of the posterior distribution, such as its moments. MCMC algorithms aim to solve this task relying on a Markov chain, $({X^{\mathcal{D}}_{n}})_{n\geqslant 1}$, which has law that converges to $\pi_{\mathcal{D}}$ asymptotically in the number of iterations $n$. Throughout the paper, we shall denote the transition kernel of the Markov chain as $P_{\mathcal{D}}:\mathrm{E}\times\mathcal{B}(\mathrm{E})\to[0,1]$, where this means ${X^{\mathcal{D}}_{n}}\sim P_{\mathcal{D}}(X^{\mathcal{D}}_{n-1},\cdot)$ and also that ${X^{\mathcal{D}}_{n}}\sim P^{n}_{\mathcal{D}}(X^{\mathcal{D}}_{0},\cdot)$, where $X^{\mathcal{D}}_{0}$ is the initial state of the chain. In particular, MCMC algorithms can be used to estimate expectations $\pi_{\mathcal{D}}(f):=\int_{\mathrm{E}}f(\theta)\pi_{\mathcal{D}}(\mathrm{d}\theta)$, for some statistics $f:\mathrm{E}\to\mathbb{R}$ thanks to associated Monte Carlo averages $\frac{1}{N}\sum_{n=1}^{N}f({X^{\mathcal{D}}_{n}})$, in the sense that $\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}f({X^{\mathcal{D}}_{n}})=\pi_{\mathcal{D}}(f)$ (Robert & Casella, 2004). There can be several possible outputs of an MCMC algorithm, such as the state of the chain at time $n$, ${X^{\mathcal{D}}_{n}}$, the entire path up to time $n$, $(X^{\mathcal{D}}_{1},\dots,{X^{\mathcal{D}}_{n}})$), or also the Monte Carlo estimator $\frac{1}{N}\sum_{n=1}^{N}f({X^{\mathcal{D}}_{n}})$. In either case, we can think of our MCMC methods as a randomised algorithm $\mathcal{A}({\mathcal{D}})$, that is a random function of the dataset. We refer to Roberts & Rosenthal (2004) for an introduction to the main concepts required to obtain (asymptotically) valid MCMC algorithms.
We refer to as randomized algorithm any function ${\mathcal{D}}\mapsto\mathcal{A}({\mathcal{D}})$, such that $\mathcal{A}({\mathcal{D}})$ is a random variable on a probability space $(\Omega,\mathbb{P},\mathcal{F})$. The framework of differential privacy (Dwork, 2006) compares the output of a randomised algorithm obtained giving two adjacent datasets as input. Various notions of adjacency between datasets can be introduced, where the most popular considers adjacent any two datasets that differ in only one entry. Below we state the definition of differential privacy.
@@
$\mathbb{P}(\mathcal{A}({\mathcal{D}})\in\mathsf{B})\leqslant e^{\varepsilon}\,\mathbb{P}(\mathcal{A}({\mathcal{D}}^{\prime})\in\mathsf{B})+\delta,$ (2)
@@
Following Definition 2.1, we say that a posterior distribution is $(\varepsilon,\delta)$-differentially private when, for all measurable sets $\mathsf{B}$, it holds that
@@
$\mathsf{D}_{\alpha}(P\rVert Q):=\frac{1}{\alpha-1}\log\mathbb{E}_{Q}\left[\left(\frac{\mathrm{d}P}{\mathrm{d}Q}\right)^{\alpha}\right].$ (4)
@@
When $\alpha\to\infty$ we recover $(\varepsilon,0)$-DP. Notice also that a $(\alpha,\varepsilon)$-Rényi-differentially private algorithm is also $(\varepsilon-\frac{1}{\alpha-1}\log\delta,\delta)$-differentially private for any $\delta\in(0,1)$ (see Proposition 3 in Mironov (2017)). Similarly to (3), we say a posterior distribution is $(\alpha,\varepsilon)$-Rényi-differentially private when $\mathsf{D}_{\alpha}(\pi_{\mathcal{D}},\pi_{{\mathcal{D}}^{\prime}})\leqslant\varepsilon$ for all adjacent datasets ${\mathcal{D}},{\mathcal{D}}^{\prime}$.
@@ 3.1 Differential privacy of a Markov chain after $n$ steps
In order to address the question above, we consider the randomised algorithm $\mathcal{A}_{s}({\mathcal{D}})=X_{n}^{\mathcal{D}}$, where $(X_{k}^{\mathcal{D}})_{k\in\mathbb{N}}$ is a Markov chain with transition kernel $P_{\mathcal{D}}$ and that is initialised from a probability distribution $\nu_{\mathcal{D}}.$ The law of $X_{n}^{\mathcal{D}}$ is denoted as $\nu_{\mathcal{D}}P_{\mathcal{D}}^{n}(\cdot):=\int\nu_{\mathcal{D}}(\mathrm{d}x)P_{\mathcal{D}}^{n}(x,\cdot)$. If $\|\cdot\|_{\mathrm{TV}}$ denotes the total variation distance (see Definition A.1), for two probability distributions $\mu_{\mathcal{D}}$, $\nu_{\mathcal{D}}$ that satisfy $\lVert\mu_{\mathcal{D}}-\nu_{\mathcal{D}}\rVert_{\mathrm{TV}}\leqslant\beta$, we have that $(\varepsilon,\delta)$-DP of $\mu_{\mathcal{D}}$ implies $(\varepsilon,\delta+\beta(e^{\varepsilon}+1))$-DP of $\mu_{\mathcal{D}}$ (see Proposition A.2 in Section A.1 for the proof of this result). In the case of MCMC algorithms, the law of $\mathcal{A}({\mathcal{D}})$ depends on the number of iterations $n,$ and so will its total variation distance to $\pi_{\mathcal{D}}.$ Nevertheless, also in our case we can apply Proposition A.2 to address the question above. In this sense we make the following assumption, which requires a bound on the total variation distance that is uniform over datasets and decreasing in the number of iterations.
@@
In particular, Assumption 3.1 requires that $\zeta$ and $\mathsf{R}$ are independent of the dataset. The function $\mathsf{R}$ is typically such that either $\lim_{m\to\infty}\mathsf{R}(m)=\underline{r}=0$, that is the when case the MCMC algorithm is asymptotically exact, or $\underline{r}>0$, that is the case when the MCMC algorithm is biased, for instance as a result of using an unadjusted discretisation scheme of a continuous time process. Under this assumption, we can obtain the following relations between the DP guarantees of $\mathcal{A}_{s}({\mathcal{D}})$ and $\pi_{\mathcal{D}}$.
@@
Consider the randomised algorithm $\mathcal{A}_{s}({\mathcal{D}})\sim\nu_{\mathcal{D}}P_{\mathcal{D}}^{n}$ and suppose Assumption 3.1 is verified. The following statements hold:
@@
The first statement is obtained applying Proposition A.2. The second statement follows applying Proposition A.2 to obtain that $\pi_{\mathcal{D}}$ is $(\varepsilon,\delta+(1+e^{\varepsilon})\zeta\,\mathsf{R}(n))$-differentially private, then taking the limit as $n\to\infty$. ∎
The first statement is similar to Proposition 3 in Wang et al. (2015). We note also that the second statement in Proposition 3.2 is still valid when the convergence bound (5) is not data-uniform.
The following result is a simple consequence of the second statement in Proposition 3.2.
Corollary 3.3.
Consider the randomised algorithm $\mathcal{A}_{s}({\mathcal{D}})\sim\nu_{\mathcal{D}}P_{\mathcal{D}}^{n}$ and suppose Assumption 3.1 is verified. Let $\varepsilon\geqslant 0$, $\delta\in[0,1)$, and $\delta^{\prime}=\delta+(1+e^{\varepsilon})\zeta\underline{r}$. If $\pi$ is not $(\varepsilon,\delta^{\prime})$-differentially private, then there exists $n\in\mathbb{N}$ such that $\mathcal{A}_{s}({\mathcal{D}})$ is not $(\varepsilon,\delta)$-differentially private.
@@
The result holds by contradiction using the second statement in Proposition 3.2. ∎
Corollary 3.3 highlights the importance of starting with a differentially private posterior distribution. Indeed, any MCMC algorithm that is asymptotically exact, i.e. such that $\underline{r}=0$, will fail to be $(\varepsilon,\delta)$-differentially private for some $n$ when the posterior itself is not $(\varepsilon,\delta)$-differentially private.
Finally, we obtain a lower bound for $\lVert\nu_{\mathcal{D}}P_{\mathcal{D}}^{n}-\pi_{\mathcal{D}}\rVert_{\mathrm{TV}}$ when $P_{{\mathcal{D}}}^{n}$ is $(\varepsilon,\delta_{n})$-differentially private, while the target distribution is not $(\varepsilon,\delta)$-differentially private for some $\delta>\delta_{n}$. The result below shows that $P^{n}_{\mathcal{D}}$ can be very far from $\pi_{\mathcal{D}}$ (in TV distance) when $\delta$ is larger than $\delta_{n}$.
@@
Let $\{P^{n}_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ be a family of Markov kernels and $\{\nu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ a family of initial distribution such that for any $n\geqslant 1$, the randomised algorithm $\mathcal{A}({\mathcal{D}})\sim\nu_{{\mathcal{D}}}P_{{\mathcal{D}}}^{n}$ is $(\varepsilon,\delta_{n})$-differentially private, and let $\{\pi_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ be a family of probability distributions that is not $(\varepsilon,\delta)$-differentially private, where $\delta>\delta_{n}$. Then, there exists a dataset ${\mathcal{D}}\in\mathcal{S}$ such that
$\lVert\nu_{{\mathcal{D}}}P_{\mathcal{D}}^{n}-\pi_{\mathcal{D}}\rVert_{\mathrm{TV}}>\frac{e^{-\varepsilon}}{1+e^{-\varepsilon}}(\delta-\delta_{n})\,.$ (6)
@@
The statement follows by an application of Proposition A.3 in Section A.2. ∎
The statement of Proposition A.3 in Section A.2 gives more insight than the statement above. Indeed, it shows that, for any pair of adjacent datasets ${\mathcal{D}},{\mathcal{D}}^{\prime}$ for which there exists a measurable set $\mathsf{B}$ such that $\pi_{\mathcal{D}}(\mathsf{B})>e^{\varepsilon}\pi_{{\mathcal{D}}^{\prime}}(\mathsf{B})+\delta$, (6) holds for either ${\mathcal{D}}$ or ${\mathcal{D}}^{\prime}$. This means that the $n$-th state of the MCMC algorithm can be far from the posterior distribution for numerous datasets. In practice, a posterior distribution $\pi$ can violate $(\varepsilon,\delta)$-DP even for large $\varepsilon$ and $\delta$ unless it is carefully designed (Dimitrakakis et al., 2017). Hence, Proposition 3.4 shows that it is essential to design $\pi$ carefully.

### v2

@@ 1 Introduction
Several works obtain DP of the one-step transitions of the MCMC algorithm by injecting extra noise in various ways, e.g. in the acceptance-rejection step of Metropolis-Hastings-type algorithms (Heikkilä et al., 2019; Yıldırım & Ermics, 2019), or leveraging the additional randomness already introduced by subsampling strategies (Wang et al., 2015; Bierkens & Duncan, 2022), optionally together with clipping of the gradient that drives the moves of the chain (Song et al., 2013; Abadi et al., 2016). Many iterations of the MCMC chain are required to approach the posterior distribution, and the typical strategy to obtain guarantees on the DP of the algorithm after some number $n$ of iterations is to use composition bounds (Abadi et al., 2016; Wang et al., 2019; Bassily et al., 2014; Ganesh et al., 2023). However, the privacy implied using composition bounds for Markov chains naturally decays with the number of steps (Kairouz et al., 2015) and as a consequence the injected noise must scale with the number of iterations in order to obtain uniform privacy bounds for the final draw. Many other works using different techniques to prove uniform-in-time privacy bounds for the final draw from a Markov chain have bad dependence on the radius of the state space or the stepsize, see (Altschuler & Talwar, 2022; Asoodeh & Diaz, 2023). These limitations were overcome in Chourasia et al. (2021), in which uniform-in-time bounds on the (Rényi) DP of gradient descent-type algorithms was obtained, however under a strong convexity assumption.
@@
In the first part (Section˜3) we consider general MCMC algorithms and establish clear connections between the DP guarantees of the posterior distribution and of a corresponding MCMC algorithm. Assuming DP of the posterior, Propositions˜3.2 and 3.10 show $(\varepsilon,\delta)$-DP for the MCMC algorithm when we either release the $n$-th state of the chain or a (noisy) Monte Carlo estimator. These result show that, under suitable assumptions, the MCMC algorithm inherits the “good" privacy properties of the posterior. Analogously, Propositions˜3.3 and 3.4 illustrate how “bad" privacy properties of the posterior affect the privacy of an MCMC algorithm. In particular, Proposition˜3.3 shows how an MCMC algorithm violates DP at a certain level $(\varepsilon,\delta)$ whenever it is sufficiently close to the posterior, assuming the posterior is not $(\varepsilon,\delta^{\prime})$-DP for some $\delta^{\prime}>\delta.$ On the other hand, Proposition˜3.4 shows that if the posterior has weaker DP guarantees than the MCMC algorithm, then the law of the MCMC chain after $n$ iterations can be far from the posterior in total variation distance. These results emphasise the importance of starting with a differentially private posterior, rather than only focusing on the one-step DP of the MCMC algorithm as often suggested in the literature. This viewpoint guarantees that the output of the MCMC algorithm can then be both private and close to the posterior.
In Section˜4 we prove privacy bounds for a class of Markov chains in a non-convex setting, using a novel approach to establish uniform-in-time $(\varepsilon,\delta)$-DP and Rényi-DP guarantees of the $n$th state of the chain, as well as bounds for the trajectory up to the $n$th state. We show in Section˜4.1 that it is possible to prove privacy guarantees via high probability bounds on the Radon-Nikodym derivative of different probability measures on the underlying probability space. In Section˜4.2 we illustrate how this strategy can be applied to the case of MCMC algorithms based on diffusions, where we rely on Girsanov’s theorem to express the Radon-Nikodym derivative. Using a careful perturbation technique, we are able to apply our approach to the case where only the final state after $n$ iterations is released, and not the entire path. In this case, we obtain uniform-in-time guarantees both for DP and Rényi-DP. Our strategy of proof also allows to obtain bounds for the entire trajectory of a variety of stochastic algorithms, improving on $(\epsilon,\delta)$-DP composition bounds and matching Rényi composition bounds. In particular, we show that the entire trajectory up to the $n$-th iteration is $(\varepsilon,\delta)$-DP, where $\varepsilon$ is $O(n+\sqrt{n\log(1/\delta)})$ and $\delta>0$ is constant, and also $(\alpha,\varepsilon)$-Rényi-DP with $\varepsilon=O(\alpha n)$ for $\alpha>0$ constant. These techniques have the additional advantage of extending easily to the continuous time case. We then focus our analysis on two classical MCMC algorithms based on the Langevin diffusion: the unadjusted Langevin algorithm and its stochastic gradient variant (see Section˜5). In doing so, we address the open problem of obtaining uniform-in-time DP guarantees in a non-convex setting.
@@ 2 Preliminaries
$\pi_{\mathcal{D}}(\mathsf{B})=\left.\int_{\mathsf{B}}L_{\mathcal{D}}(\theta)\lambda(\mathrm{d}\theta)\middle/{Z_{\mathcal{D}}}\right.\;,\quad\mathsf{B}\in\mathcal{B}(\mathrm{E})\;,$ (2)
where $(\mathrm{E},\mathcal{B}(\mathrm{E}))$ is a Borel space, $L_{\mathcal{D}}(\theta)$ is the likelihood associated with $\mathcal{D}$}, $\lambda$ is a prior distribution for the parameter, and $Z_{\mathcal{D}}=\int_{\mathrm{E}}L_{\mathcal{D}}(\theta)\eta(\mathrm{d}\theta)$. In order to make practical use of the Bayesian framework, it is then essential to have access to key statistics of the posterior distribution, such as its moments. MCMC algorithms aim to solve this task relying on a Markov chain, $({X^{\mathcal{D}}_{n}})_{n\geqslant 1}$, which has law that converges to $\pi_{\mathcal{D}}$ asymptotically in the number of iterations $n$. Throughout the paper, we shall denote the transition kernel of the Markov chain as $P_{\mathcal{D}}:\mathrm{E}\times\mathcal{B}(\mathrm{E})\to[0,1]$, where this means ${X^{\mathcal{D}}_{n}}\sim P_{\mathcal{D}}(X^{\mathcal{D}}_{n-1},\cdot)$ and also that ${X^{\mathcal{D}}_{n}}\sim P^{n}_{\mathcal{D}}(X^{\mathcal{D}}_{0},\cdot)$, where $X^{\mathcal{D}}_{0}$ is the initial state of the chain. In particular, MCMC algorithms can be used to estimate expectations $\pi_{\mathcal{D}}(f):=\int_{\mathrm{E}}f(\theta)\pi_{\mathcal{D}}(\mathrm{d}\theta)$, for some statistics $f:\mathrm{E}\to\mathbb{R}$ thanks to associated Monte Carlo averages $\frac{1}{N}\sum_{n=1}^{N}f({X^{\mathcal{D}}_{n}})$, in the sense that $\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}f({X^{\mathcal{D}}_{n}})=\pi_{\mathcal{D}}(f)$ (Robert & Casella, 2004). There can be several possible outputs of an MCMC algorithm, such as $X^{\mathcal{D}}_{n}$}, the state of the chain at time $n$, or $(X^{\mathcal{D}}_{1},\dots,{X^{\mathcal{D}}_{n}})$), the entire path up to time $n$, or also the Monte Carlo estimator $\frac{1}{N}\sum_{n=1}^{N}f({X^{\mathcal{D}}_{n}})$. In either case, we can think of our MCMC methods as a randomised algorithm $\mathcal{A}({\mathcal{D}})$, that is a random function of the dataset. We refer to Roberts & Rosenthal (2004) for an introduction to the main concepts required to obtain (asymptotically) valid MCMC algorithms.
We call randomised algorithm any function ${\mathcal{D}}\mapsto\mathcal{A}({\mathcal{D}})$, such that $\mathcal{A}({\mathcal{D}})$ is a random variable on a probability space $(\Omega,\mathbb{P},\mathcal{F})$. We denote the law of a randomised algorithm $\mathcal{A}$ as $P_{\mathcal{A}({\mathcal{D}})}$, that is defined for any measurable set $\mathsf{B}$ as
$P_{\mathcal{A}({\mathcal{D}})}(\mathsf{B}):=\mathbb{P}(\mathcal{A}({\mathcal{D}})\in\mathsf{B}).$ (3)
The framework of differential privacy (Dwork, 2006) compares the output of a randomised algorithm obtained giving two adjacent datasets as input. Various notions of adjacency between datasets can be introduced, where the most popular considers adjacent any two datasets that differ in only one entry. Below we state the definition of differential privacy.
@@
$P_{\mathcal{A}({\mathcal{D}})}(\mathsf{B})\leqslant e^{\varepsilon}\,P_{\mathcal{A}({\mathcal{D}}^{\prime})}(\mathsf{B})+\delta,$ (4)
@@
Following Definition˜2.1, we say that a posterior distribution is $(\varepsilon,\delta)$-differentially private when, for all measurable sets $\mathsf{B}$, it holds that
@@
$\mathsf{D}_{\alpha}(P\|Q):=\frac{1}{\alpha-1}\log\biggr(\int_{\Omega}\left(\frac{\mathrm{d}P}{\mathrm{d}Q}\right)^{\alpha}dQ\biggr).$ (6)
@@
When $\alpha\to\infty$ we recover $(\varepsilon,0)$-DP. Notice also that a $(\alpha,\varepsilon)$-Rényi-differentially private algorithm is also $(\varepsilon-\frac{1}{\alpha-1}\log\delta,\delta)$-differentially private for any $\delta\in(0,1)$ (see Proposition 3 in Mironov (2017)). Similarly to (5), we say a posterior distribution is $(\alpha,\varepsilon)$-Rényi-differentially private when $\mathsf{D}_{\alpha}(\pi_{\mathcal{D}}\|\pi_{{\mathcal{D}}^{\prime}})\leqslant\varepsilon$ for all adjacent datasets ${\mathcal{D}},{\mathcal{D}}^{\prime}$.
@@ 3.1 Differential privacy of a Markov chain after $n$ steps
In order to address the question above, we consider the randomised algorithm $\mathcal{A}_{s}({\mathcal{D}})=X_{n}^{\mathcal{D}}$, where $(X_{k}^{\mathcal{D}})_{k\in\mathbb{N}}$ is a Markov chain with transition kernel $P_{\mathcal{D}}$ and that is initialised from a probability distribution $\nu_{\mathcal{D}}.$ The law of $X_{n}^{\mathcal{D}}$ is denoted as $\nu_{\mathcal{D}}P_{\mathcal{D}}^{n}(\cdot):=\int\nu_{\mathcal{D}}(\mathrm{d}x)P_{\mathcal{D}}^{n}(x,\cdot)$. If $\|\cdot\|_{\mathrm{TV}}$ denotes the total variation distance (see Definition˜A.1), for two families of probability distributions $\mu_{\mathcal{D}}$, $\nu_{\mathcal{D}}$ that satisfy $\lVert\mu_{\mathcal{D}}-\nu_{\mathcal{D}}\rVert_{\mathrm{TV}}\leqslant\beta$, we have that $(\varepsilon,\delta)$-DP of $\mu_{\mathcal{D}}$ implies $(\varepsilon,\delta+\beta(e^{\varepsilon}+1))$-DP of $\nu_{\mathcal{D}}$ (see Proposition˜A.2 in Section˜A.1 for the proof of this result). In the case of MCMC algorithms, the law of $\mathcal{A}({\mathcal{D}})$ depends on the number of iterations $n,$ and so will its total variation distance to $\pi_{\mathcal{D}}.$ Nevertheless, also in our case we can apply Proposition˜A.2 to address the question above. In this sense we make the following assumption, which requires a bound on the total variation distance that is uniform over datasets and decreasing in the number of iterations.
@@
In particular, Assumption˜3.1 requires that $\zeta$ and $\mathsf{R}$ are independent of the dataset. The function $\mathsf{R}$ is typically such that either $\lim_{m\to\infty}\mathsf{R}(m)=\underline{r}=0$, that is the when case the MCMC algorithm is asymptotically exact, or $\underline{r}>0$, that is the case when the MCMC algorithm is biased, for instance as a result of using an unadjusted discretisation scheme of a continuous time process. Under this assumption, we can obtain the following relations between the DP guarantees of $\mathcal{A}_{s}({\mathcal{D}})$ and $\pi_{\mathcal{D}}$.
@@
Consider the randomised algorithm $\mathcal{A}_{s}({\mathcal{D}})\sim\nu_{\mathcal{D}}P_{\mathcal{D}}^{n}$ and suppose Assumption˜3.1 is verified. The following statements hold:
@@
The first statement is obtained applying Proposition˜A.2. The second statement follows applying Proposition˜A.2 to obtain that $\pi_{\mathcal{D}}$ is $(\varepsilon,\delta+(1+e^{\varepsilon})\zeta\,\mathsf{R}(n))$-differentially private, then taking the limit as $n\to\infty$. ∎
The first statement is similar to Proposition 3 in Wang et al. (2015). We note also that the second statement in Proposition˜3.2 is still valid when the convergence bound (7) is not data-uniform.
A simple corollary of Proposition˜3.2-(ii) is that any MCMC algorithm that is asymptotically exact, i.e. such that $\underline{r}=0$, will fail to be $(\varepsilon,\delta)$-differentially private for some number of iterations $n$ when the posterior itself is not $(\varepsilon,\delta)$-differentially private. The following proposition gives a quantitative result in this direction.
Proposition 3.3.
Suppose Assumption˜3.1 is verified. Let $\varepsilon\geqslant 0$, $\delta\in[0,1)$, and $\delta^{\prime}\in(\delta,1)$ and suppose $\pi_{\mathcal{D}}$ is not $(\varepsilon,\delta^{\prime})$-differentially private. Let
$n^{*}=\inf\{n\in\mathbb{N}:\,(1+e^{\varepsilon})\zeta\mathsf{R}(n)\leqslant\delta^{\prime}-\delta\}.$
Then, $\mathcal{A}_{s}({\mathcal{D}})\sim\nu_{\mathcal{D}}P_{\mathcal{D}}^{n}$ is not $(\varepsilon,\delta)$-differentially private for all $n\geqslant n^{*}$.
@@
The result follows applying Proposition˜A.3, that can be found in Section˜A.2. ∎
Proposition˜3.3 gives that $\mathcal{A}_{s}({\mathcal{D}})\sim\nu_{\mathcal{D}}P_{\mathcal{D}}^{n}$ is not $(\varepsilon,\delta)$-differentially private whenever it is sufficiently close to $\pi_{\mathcal{D}}$, assuming $\pi_{\mathcal{D}}$ is not $(\varepsilon,\delta^{\prime})$-differentially private for some $\delta^{\prime}>\delta.$
Finally, we show that $\nu_{\mathcal{D}}P_{\mathcal{D}}^{n}$ can be far from $\pi_{\mathcal{D}}$ when $\nu_{\mathcal{D}}P_{{\mathcal{D}}}^{n}$ is $(\varepsilon,\delta_{n})$-differentially private, while the posterior distribution violates a weaker DP guarantee. In particular, we say that the posterior distribution is not $(\varepsilon,\delta)$-differentially private when there exist adjacent datasets ${\mathcal{D}},{\mathcal{D}}^{\prime}\in\mathcal{S}$ and a measurable set $\mathsf{B}$ such that $\pi_{\mathcal{D}}(\mathsf{B})>e^{\varepsilon}\pi_{{\mathcal{D}}^{\prime}}(\mathsf{B})+\delta$.
@@
Consider the randomised algorithm $\mathcal{A}({\mathcal{D}})\sim\nu_{{\mathcal{D}}}P_{{\mathcal{D}}}^{n}$ and assume it is $(\varepsilon,\bar{\delta})$-differentially private. Let $\{\pi_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ be a family of posterior distributions that is not $(\varepsilon,\delta)$-differentially private for some $\delta>\bar{\delta}$. Then, there exists a dataset ${\mathcal{D}}\in\mathcal{S}$ such that
$\lVert\nu_{{\mathcal{D}}}P_{\mathcal{D}}^{n}-\pi_{\mathcal{D}}\rVert_{\mathrm{TV}}>\frac{e^{-\varepsilon}}{1+e^{-\varepsilon}}(\delta-\bar{\delta})\,.$ (8)
@@
The statement follows by an application of Proposition˜A.4 in Section˜A.3. ∎
The statement of Proposition˜A.4 in Section˜A.3 gives more insight than the statement above. Indeed, it shows that, for any pair of adjacent datasets ${\mathcal{D}},{\mathcal{D}}^{\prime}$ for which there exists a measurable set $\mathsf{B}$ such that $\pi_{\mathcal{D}}(\mathsf{B})>e^{\varepsilon}\pi_{{\mathcal{D}}^{\prime}}(\mathsf{B})+\delta$, (8) holds for either $\mathcal{D}$} or $\mathcal{D}$$^{\prime}$. This means that the $n$-th state of the MCMC algorithm can be far from the posterior distribution for numerous datasets. In practice, a posterior distribution $\pi$ can violate $(\varepsilon,\delta)$-DP even for large $\varepsilon$ and $\delta$ unless it is carefully designed (Dimitrakakis et al., 2017). Hence, Proposition˜3.4 shows that it is essential to design $\pi$ carefully.